Birdseed costs ​$0.56 a pound and sunflower seeds cost ​$0.86 a pound. Angela​ Leinenbachs' pet store wishes to make a 40​-pound mixture of birdseed and sunflower seeds that sells for ​$0.75 per a pound. How many pounds of each type of seed should she​ use?
Answer by Alan3354(69443)

Birdseed costs ​$0.56 a pound and sunflower seeds cost ​$0.86 a pound. Angela​ Leinenbachs' pet store wishes to make a 40​-pound mixture of birdseed and sunflower seeds that sells for ​$0.75 per a pound. How many pounds of each type of seed should she​ use?
==================
b + s = 40 --- total mix
56b + 86s = 40*75 ---- cost in cents